# Analyzing the Usage of Donation Platforms for PyPI Libraries

### v1

@@ 6. Threats to Validity
Internal validity: No threat, as we focus on describing ecosystem characteristics rather than examining causal relationships.
External validity: We concentrate on PyPI as the ecosystem and GitHub as the code management system (CMS). While the findings might not apply universally to other ecosystems and CMSs, the research methodology applied can be adapted to any publicly accessible ecosystem where libraries can incorporate links, like those to donation platforms and code repositories, within their metadata. We did not consider alternative CMSs, as previous research showed that PyPI libraries strongly prefer linking to GitHub repositories over alternatives like GitLab (Tsakpinis and Pretschner, 2024).
Reliability: Although we gathered all data from publicly accessible sources, which theoretically supports reproducibility, a threat to reliability arises from the dynamic nature of the PyPI ecosystem. The data represent a specific snapshot from the collection day, which conflicts with reproducibility since PyPI does not offer any time travel functionality. To facilitate full reproducibility, the collected data and analysis code are made available on figshare (Anonymous, 2025). Another risk comes from the frequent introduction or modification of libraries, which could lead to gaps or outdated information in our dataset. To mitigate this, the study should be repeated regularly.
@@ 7. Conclusion and Future Work
The paper analyzes how maintainers of PyPI libraries utilize donation platforms on PyPI and GitHub. These platforms play a crucial role in providing infrastructure for financial support, which is vital for the sustainability of OSS projects, as a lack of funding opportunities can lead to reduced maintenance activities, ultimately increasing security risks. We analyzed the PyPI ecosystem from a library and dependency chain perspective, offering a comprehensive overview of how donation platforms are utilized. From a library perspective, we discovered that donation platform links are mostly missing on PyPI project pages, with a clear tendency to list them on GitHub repositories instead. GitHub Sponsors stands out as the primary donation platform across PyPI and GitHub. From a dependency chain perspective, we discovered that a significant proportion of libraries link to donation platforms — up to 72.8% in direct dependencies and 76.4% in transitive ones — demonstrating that most dependencies within the PyPI ecosystem are eligible to receive funding from end users or industry initiatives.

### v3

@@ 6. Threats to Validity
Internal validity: Threats may stem from data‑handling errors (e.g., mis‑tagged URLs, missed dependencies). We mitigated them by rerunning the data collection and spot‑checking random samples.
External validity: We concentrate on PyPI as the ecosystem and GitHub as the code management system (CMS). While the findings might not apply universally to other ecosystems and CMSs, the research methodology applied can be adapted to any publicly accessible ecosystem where libraries can incorporate links, like those to donation platforms and code repositories, within their metadata. We did not consider alternative CMSs, as previous research showed that PyPI libraries strongly prefer linking to GitHub repositories over alternatives like GitLab or BitBucket (Tsakpinis and Pretschner, 2024).
Reliability: Although we gathered all data from publicly accessible sources, which theoretically supports reproducibility, a threat to reliability arises from the dynamic nature of the PyPI ecosystem. The data represent a specific snapshot from the collection day, which conflicts with reproducibility since PyPI does not offer any time travel functionality. To facilitate full reproducibility, the collected data and analysis code are made available on figshare (tsa, 2025). Another risk comes from the frequent introduction or modification of libraries, which could lead to gaps or outdated information in our dataset. To mitigate this, the study should be repeated regularly.
@@ 7. Conclusion and Future Work
The paper analyzes how maintainers of PyPI libraries utilize donation platforms on PyPI and GitHub. These platforms play a crucial role in providing infrastructure for financial support, which is vital for the sustainability of OSS projects, as a lack of funding opportunities can lead to reduced maintenance activities, ultimately increasing security risks. We analyzed the PyPI ecosystem from a library and dependency chain perspective, offering a comprehensive overview of how donation platforms are utilized. From a library perspective, we discovered that donation platform links are mostly missing on PyPI project pages, with a clear tendency to list them on GitHub repositories instead. GitHub Sponsors stands out as the primary donation platform across PyPI and GitHub. From a dependency chain perspective, we discovered that a significant proportion of libraries link to donation platforms — up to 72.8% in direct dependencies and 76.4% in transitive ones — demonstrating that most dependencies within the PyPI ecosystem are eligible to receive funding from end users and industry initiatives.